# Phaedrus: Predicting Dynamic Application Behavior with Lightweight Generative Models and LLMs

### v3

@@ 1. Introduction
Application profiling and execution tracing are crucial in the software development process. Popular performance enhancement techniques such as profile-guided optimization (PGO) (Radigan et al., 1995; Grove et al., 1995; Gupta et al., 1997) involve generating execution traces for applications and then performing optimization decisions specific to the collected profile, such as code transformation (Chang et al., 1991; Visochan et al., 2022; Sydow et al., 2020; Joshi et al., 2004), data layout restructuring and reorganization (Pettis and Hansen, 1990). Application profiles and traces are also used to accomplish a wide variety of tasks such as estimating resource usage for scheduling and resource allocation (Agelastos et al., 2015; Song et al., 2009; Viswanathan et al., 2011; Rashti et al., 2015; Agelastos et al., 2016; Kestor et al., 2013), memory management for scratchpads, caches, and other memory hierarchies (Cho et al., 2009; Ansari et al., 2009; Bruno and Ferreira, 2017; Agarwal and Wenisch, 2017; Xu et al., 2024), memory simulation (Uhlig and Mudge, 1997, 2001; Rico et al., 2011; Hassan et al., 2007) and debugging (Arnold et al., 2007; Nagarajan et al., 2012; Ning and Zhang, 2017, 2018; Verma et al., 2020). Such profile-driven approaches have been successfully utilized and deployed for over two decades.
Most profile-driven tasks rely on the users to manually select profiles representative of the application’s intended execution scenario. This ensures the usefulness of the profile analysis and optimization decisions during the application’s runtime across different inputs. However, recent work (Chatterjee et al., 2022; Mururu et al., 2023; Ahmed et al., 2017) has shown that most modern applications are complex and exhibit a high amount of input dependence, resulting in variable application behavior across different execution instances on various input data sets. This variant behavior is exhibited in all program components, including program branches, loops, and function calls. As a result, the application profile collected during a particular execution instance on a given input data set can be entirely different from another execution instance collected using a different input data set. A naive approach to tackle this problem would be to profile the application on multiple inputs - in practice, however, this ‘excessive’ profiling approach does not scale for workloads whose input domain is not numerable, such as graph applications. In addition, profile mixing can also be problematic for specific path-based optimizations that optimize along hot trace paths, pushing overheads to off-trace paths.
@@
Phaedrus Framework: In a nutshell, the idea of this work is to utilize generative models to predict the behavior of dynamic applications. The predictions from these models are then further used to perform profile-guided optimization. Towards this, the first new concept introduced in this work is application behavior synthesis, which attempts to reason about the dynamic program behavior without execution or collected profiles with the aid of Large-Language Models (LLMs). Owing to their remarkable successes in every domain, from natural language processing to software engineering, LLMs have been extensively used for generating and optimizing programs (Ugare et al., 2024; Gu, 2023; Lin et al., 2024; Liu et al., 2024). However, the primary challenge that prevents the adoption of LLM in place of the mainstream compilation pipeline, is the absence of correctness guarantees (Zhong and Wang, 2024; Spiess et al., 2024). Secondly, LLMs, being generative models at their core, cannot inherently perform static program analysis, which can enhance its decision-making capacities. Thus, naively querying the LLM for most code reasoning tasks fails spectacularly (as shown in fig. 1).
@@
Compiler-Assisted LLM Reasoning: The first contribution of this work is the utilization of Large-Language Models, supplemented by the knowledge of static program properties obtained by compiler analysis. The Dynamis framework operates in three broad phases: first, filtering the wide corpus of LLM knowledge through domain knowledge deduction3.1); second, supplementing this understanding with program structural artifacts such as the universe of functions and call graphs (§3.2); and finally, fortifying the LLM’s reasoning with static compiler analysis3.3). While recent efforts have focused on using LLMs to generate optimized code (Gao et al., 2024; Qiu et al., 2024; Rosas et al., 2024), such code often requires extensive verification (Liu et al., 2023b; Tang, 2024; Ngassom et al., 2024). Unlike these approaches, Dynamis leverages LLMs only to identify input-specific profile-driven optimization opportunities, followed by traditional code generation techniques to elicit performance.
However, a downside of using LLMs for determining input-specific program behavior, is resource requirements needed for successful deployment and inference. For instance, deploying a state-of-art LLM such as LLaMA-2-70B with 70 billion parameters for inference, can take multiple NVIDIA A100 GPUs (Zhou et al., 2024a). Secondly, LLMs often have a fixed context window that determines the amount of tokens that it can consider to answer a certain query (8192 tokens for GPT-4). Large applications with hundreds of KLOC can easily exceed this limit, and lead to undesirable results. Most importantly, with the set of prompts used in the current work, we observe that LLMs are unable to reason about the dynamic sequence of function calls satisfactorily (§5.6). In the absence of accurate function call sequence, most path-based optimizations and execution forensics cannot be achieved.
@@
As a lightweight alternative to the LLM-based approach, in this work, we leverage Whole Program Paths (WPP) based function profile to make decisions based on the program’s runtime execution behavior. Whole Program Path Profiles (Larus, 1999; Ball and Larus, 2000; Zhang and Gupta, 2001; Tallam et al., 2005) captures the entire dynamic control flow of an application during its execution, and are superior in providing insight about the dynamic execution behavior than popular sampling-based trace collection methods (Martonosi et al., 1993; Kessler et al., 1994; Novillo, 2014; Liu et al., 2016). However, the biggest challenge that prevents the adoption of WPP Profiles for profile-driven tasks in modern workloads is its intractability with respect to data size and volume (§5.5). Although compression of WPP profiles has been explored extensively in compiler literature (Larus, 1999; Zhang and Gupta, 2001; Burtscher and Jeeradit, 2003; Zhang and Gupta, 2002; Law and Rothermel, 2003; Milenkovic and Milenkovic, 2003; Zhang and Gupta, 2004; Lin et al., 2007), unfortunately such compression schemes can only be applied to a given WPP profile obtained from a single input data set at a time and cannot be used to get a generalized representation of profiles across different inputs (§2.3).
@@
The experimental evaluation (§5) of Phaedrus is done on the SPEC CPU 2017 benchmark suite (Bucek et al., 2018). Our evaluation focuses on Dynamis and application behavior synthesis, we evaluate the prediction of hot functions and the coverage of several function calls (§5.1). We also show that Phaedrus can achieve application binary size reduction of upto 65% (13.68% on average) and performance improvement of upto 13.34% (2.8% on average). For Morpheus and application profile generalization, our evaluation focuses on the order of compression for WPP based function profiles (upto $10^{7}\times$ reduction), prediction of hot functions and its execution coverage and training overheads (§5.5). To the best of our knowledge, this is the first work to explore dynamic application behavior across different execution scenarios and leverages compiler-assisted lightweight deep learning models and LLMs for predicting new program behaviors for profile-guided tasks. .
@@ 2.1. Can LLMs reason about the dynamic program behavior?
Large Language Models (LLMs) exhibit remarkable capabilities in terms of code understanding (Fang et al., 2024; Nam et al., 2024; Wang et al., 2023) and analysis (Fang et al., 2024; Ma et al., 2024b), making them valuable tools for developers. One of their primary strengths lies in understanding code and describing behavior. LLMs can analyze and explain the logic of code snippets, breaking down complex functions into more comprehensible explanations for developers. They are also capable of predicting potential runtime behavior (Chen et al., 2024) based on static code analysis, such as identifying loops, conditionals, and function calls that may impact the execution flow. By simulating execution paths leveraging their neural models, LLMs can infer what might happen under certain conditions, offering valuable insight into code behavior without running it. Furthermore, LLMs excel in reasoning about high-level code behavior (Ding et al., 2024) by identifying concepts like state changes (e.g., variable updates or branching behavior) and evaluating algorithmic implications, such as performance considerations and potential impacts of code modifications. This makes them useful for predicting time complexity, spotting inefficiencies, and suggesting areas for optimization (Yang et al., 2024b; Zhou et al., 2024b).
@@
However, LLMs have limitations that stem from their inability to perform deep dynamic analysis. Because they do not execute code (Xue et al., 2024; Yang et al., 2024a), they cannot analyze real-time state changes, such as tracking memory usage during execution or handling complex input/output scenarios dynamically. Their pre-trained nature also makes it challenging to reason about concurrency and parallelism, which involves understanding multi-threaded behavior, race conditions, or potential deadlocks. This type of analysis (Sun et al., 2024; Chen et al., 2024; Krishna et al., 2024) requires direct code execution and monitoring of parallel processes, which LLMs cannot achieve. Furthermore, LLMs lack the capability to handle real-time data changes or interact with code based on live data, as they primarily operate on static representations and their training knowledge. This limitation prevents them from fully simulating execution environments or adapting to live input changes, highlighting an important gap compared to runtime analysis tools.
Although LLMs are inherently limited in performing extensive dynamic analysis, particularly in complex scenarios involving real-time state changes, concurrency, and live data interactions, our framework demonstrates a path to extend their utility through strategic assistance. By combining static analysis techniques with the logical reasoning (Parmar et al., 2024; Li et al., 2024) and domain knowledge that LLMs inherently possess, we can bridge this gap effectively.
@@ 2.2. Compiler-Assisted LLM Reasoning: Combining World Knowledge with Static Analysis
When these compiler generated static artifacts are paired with an LLM’s capabilities in reasoning (Sun et al., 2024; Chen et al., 2024; Krishna et al., 2024) about control flow, algorithmic complexity, and domain-specific behavior, a comprehensive analysis of code behavior emerges. The LLM can use static analysis insights to simulate runtime behaviors, assess interactions, and predict the impact of changes more effectively. For instance, a call graph might reveal complex function dependencies, while the LLM can interpret how these dependencies influence runtime performance or suggest optimizations. However, performing such analysis within the LLM itself can be computationally expensive, shifting focus from estimating hot functions to more generalized codebase analysis. Since it is an interactive tool, LLMs typically budget a certain time interval for getting the response back to the user and within that budget different phases within the LLM must fit. To offload the burden of analysis and to utilize that time for some other inferencing phases, we assist the LLM with logical cues such as the set of all functions used within the benchmark, function call graphs, and some hotness indicators such as whether or not called within loops etc. This approach reduces the model’s search space, enabling more efficient and targeted inference by emphasizing functions that are statistically and structurally significant. This combination of static analysis and probabilistic guidance empowers the LLM to deliver more accurate and efficient code insights, bridging gaps in its inherent capabilities (Hadi et al., 2024).
@@ 2.3. Alternative Approach: Profile Generalization & Generative Models
In order to utilize this WPP based function profile data for profile generalization and accomplishing profile-driven tasks, we train generative models on WPP based function profile data, which predicts the upcoming tokens (function calls) based on a prefix of tokens, and simultaneously adjusts to other execution scenarios (different inputs). Deep-learning generative models such as Recurrent Neural Networks (RNNs) can estimate the probability of the next token (at timestep $t$) in a sequence, as function of the previous token (at timestep t1t-1) and an internal short-term memory (which propagates information across the sequence). The utility of RNNs stems from the fact that the output token in the previous execution timestep acts as the input token for the current execution timestep, which simplifies the process of querying and inferencing the model. Popular generative models such as GANs (Goodfellow et al., 2020), Diffusion Models (Yang et al., 2023) etc, are extremely compute-heavy ($\sim 10^{20}$ FLOPS per epoch), and require extensive resources for training (100s of GPU hours). In contrast, we show that simple RNNs with one hidden layer (vanilla RNNs) are capable of learning the dynamic function call chains (§5.5).
@@ 3.1. Phase 1: Filtering Search Space with Domain Knowledge Deduction
Our initial experiments showed that once the application domain is established, a direct correlation can be established between algorithm-specific cost models, and candidate hotspot functions. For instance, 510.parest_r510.parest\_r from SPEC2k17 (Bucek et al., 2018) implements a finite element solver for biomedical imaging (Listing 1). Based on this domain knowledge, the LLM can immediately deduce that quadrature evaluations, gradient computations, and scatter operations are computationally intensive. This leads to the isolation of s_val, s_grad, and cpy_matrix as potential hotspots, sidestepping the high-level assemble_matrix wrapper. Similarly, 505.mcf_r505.mcf\_r from the same benchmark suite adopts a primal network simplex algorithm to solve the minimum cost flow problem (Listing 2). This deduction enlightened the LLM to attribute runtime to the domain specific constructs such as pricing (cost_compare), sorting (spec_qsort), and implicit-arc maintenance (bea_is_dual_infeasible) functions that dominate execution at scale, instead of the top-level driver primal_net_simplex. We observed a similar pattern in other application domains, such as molecular dynamics (neighbor list construction and electrostatics kernels), fluid mechanics (collide-and-stream sweeps), and image processing (convolution, and resampling), where domain knowledge consistently filtered out high-level drivers in favor of the computational kernels.
@@
Listing 1: Finite Element Solver code snippet from 510.parest_r in SPEC2017 (Bucek et al., 2018). While the wrapper dominates inclusive time, the true hotspots are the inner quadrature and scatter kernels, based on domain knowledge.
@@
Listing 2: Primal Simplex code snippet from solving minimum cost flow in 505.mcf_r in SPEC2017 (Bucek et al., 2018). Despite enclosing the optimization loop, the runtime is dominated by the inner pricing and sorting routines

### v4

@@ 1. Introduction
Application profiling and execution tracing are crucial in the software development process. Popular performance enhancement techniques such as profile-guided optimization (PGO) (radigan1995integer; grove1995profile; gupta1997resource) involve generating execution traces for applications and then performing optimization decisions specific to the collected profile, such as code transformation (chang1991using; visochan2022method; sydow2020towards; joshi2004targeted), data layout restructuring and reorganization (pettis1990profile). Application profiles and traces are also used to accomplish a wide variety of tasks such as estimating resource usage for scheduling and resource allocation (agelastos2015toward; song2009energy; viswanathan2011energy; rashti2015wattprof; agelastos2016continuous; kestor2013enabling), memory management for scratchpads, caches, and other memory hierarchies (cho2009adaptive; ansari2009profiling; bruno2017polm2; agarwal2017thermostat; xu2024prompt), memory simulation (uhlig1997trace; uhlig2001trace; rico2011trace; hassan2007synthetic) and debugging (arnold2007stack; nagarajan2012system; ning2017ninja; ning2018hardware; verma2020interactive). Such profile-driven approaches have been successfully utilized and deployed for over two decades.
Most profile-driven tasks rely on the users to manually select profiles representative of the application’s intended execution scenario. This ensures the usefulness of the profile analysis and optimization decisions during the application’s runtime across different inputs. However, recent work (chatterjee2022cas; mururu2023beacons; ahmed2017leveraging) has shown that most modern applications are complex and exhibit a high amount of input dependence, resulting in variable application behavior across different execution instances on various input data sets. This variant behavior is exhibited in all program components, including program branches, loops, and function calls. As a result, the application profile collected during a particular execution instance on a given input data set can be entirely different from another execution instance collected using a different input data set. A naive approach to tackle this problem would be to profile the application on multiple inputs - in practice, however, this ‘excessive’ profiling approach does not scale for workloads whose input domain is not numerable, such as graph applications. In addition, profile mixing can also be problematic for specific path-based optimizations that optimize along hot trace paths, pushing overheads to off-trace paths.
@@
Phaedrus Framework: In a nutshell, the idea of this work is to utilize generative models to predict the behavior of dynamic applications. The predictions from these models are then further used to perform profile-guided optimization. Towards this, the first new concept introduced in this work is application behavior synthesis, which attempts to reason about the dynamic program behavior without execution or collected profiles with the aid of Large-Language Models (LLMs). Owing to their remarkable successes in every domain, from natural language processing to software engineering, LLMs have been extensively used for generating and optimizing programs (ugare2024improving; gu2023llm; lin2024llm; liu2024exploring). However, the primary challenge that prevents the adoption of LLM in place of the mainstream compilation pipeline, is the absence of correctness guarantees (zhong2024chatgptreplacestackoverflowstudy; spiess2024calibrationcorrectnesslanguagemodels). Secondly, LLMs, being generative models at their core, cannot inherently perform static program analysis, which can enhance its decision-making capacities. Thus, naively querying the LLM for most code reasoning tasks fails spectacularly (as shown in fig. 1).
@@
Compiler-Assisted LLM Reasoning: The first contribution of this work is the utilization of Large-Language Models, supplemented by the knowledge of static program properties obtained by compiler analysis. The Dynamis framework operates in three broad phases: first, filtering the wide corpus of LLM knowledge through domain knowledge deduction3.1); second, supplementing this understanding with program structural artifacts such as the universe of functions and call graphs (§3.2); and finally, fortifying the LLM’s reasoning with static compiler analysis3.3). While recent efforts have focused on using LLMs to generate optimized code (gao2024searchbasedllmscodeoptimization; qiu2024efficientllmgeneratedcoderigorous; rosas2024aioptimizecodecomparative), such code often requires extensive verification (liu2023; tang2024; ngassom2024). Unlike these approaches, Dynamis leverages LLMs only to identify input-specific profile-driven optimization opportunities, followed by traditional code generation techniques to elicit performance.
However, a downside of using LLMs for determining input-specific program behavior, is resource requirements needed for successful deployment and inference. For instance, deploying a state-of-art LLM such as LLaMA-2-70B with 70 billion parameters for inference, can take multiple NVIDIA A100 GPUs (zhou2024survey). Secondly, LLMs often have a fixed context window that determines the amount of tokens that it can consider to answer a certain query (8192 tokens for GPT-4). Large applications with hundreds of KLOC can easily exceed this limit, and lead to undesirable results. Most importantly, with the set of prompts used in the current work, we observe that LLMs are unable to reason about the dynamic sequence of function calls satisfactorily (§5.6). In the absence of accurate function call sequence, most path-based optimizations and execution forensics cannot be achieved.
@@
As a lightweight alternative to the LLM-based approach, in this work, we leverage Whole Program Paths (WPP) based function profile to make decisions based on the program’s runtime execution behavior. Whole Program Path Profiles (larus1999whole; ball2000using; zhang2001timestamped; tallam2005extended) captures the entire dynamic control flow of an application during its execution, and are superior in providing insight about the dynamic execution behavior than popular sampling-based trace collection methods (martonosi1993effectiveness; kessler1994comparison; novillo2014samplepgo; liu2016sample). However, the biggest challenge that prevents the adoption of WPP Profiles for profile-driven tasks in modern workloads is its intractability with respect to data size and volume (§5.5). Although compression of WPP profiles has been explored extensively in compiler literature (larus1999whole; zhang2001timestamped; burtscher2003compressing; zhang2002path; law2003whole; milenkovic2003stream; zhang2004whole; lin2007design), unfortunately such compression schemes can only be applied to a given WPP profile obtained from a single input data set at a time and cannot be used to get a generalized representation of profiles across different inputs (§2.3).
@@
The experimental evaluation (§5) of Phaedrus is done on the SPEC CPU 2017 benchmark suite (bucek2018spec). Our evaluation focuses on Dynamis and application behavior synthesis, we evaluate the prediction of hot functions and the coverage of several function calls (§5.1). We also show that Phaedrus can achieve application binary size reduction of upto 65% (13.68% on average) and performance improvement of upto 13.34% (2.8% on average). For Morpheus and application profile generalization, our evaluation focuses on the order of compression for WPP based function profiles (upto $10^{7}\times$ reduction), prediction of hot functions and its execution coverage and training overheads (§5.5). To the best of our knowledge, this is the first work to explore dynamic application behavior across different execution scenarios and leverages compiler-assisted lightweight deep learning models and LLMs for predicting new program behaviors for profile-guided tasks. .
@@ 2.1. Can LLMs reason about the dynamic program behavior?
Large Language Models (LLMs) exhibit remarkable capabilities in terms of code understanding (fang2024largelanguagemodelscode; nam2024usingllmhelpcode; wang2023codet5opencodelarge) and analysis (fang2024largelanguagemodelscode; ma2024lmsunderstandingcodesyntax), making them valuable tools for developers. One of their primary strengths lies in understanding code and describing behavior. LLMs can analyze and explain the logic of code snippets, breaking down complex functions into more comprehensible explanations for developers. They are also capable of predicting potential runtime behavior (chen2024reasoningruntimebehaviorprogram) based on static code analysis, such as identifying loops, conditionals, and function calls that may impact the execution flow. By simulating execution paths leveraging their neural models, LLMs can infer what might happen under certain conditions, offering valuable insight into code behavior without running it. Furthermore, LLMs excel in reasoning about high-level code behavior (ding2024reasoningplanningllm) by identifying concepts like state changes (e.g., variable updates or branching behavior) and evaluating algorithmic implications, such as performance considerations and potential impacts of code modifications. This makes them useful for predicting time complexity, spotting inefficiencies, and suggesting areas for optimization (yang2024large; zhou2024surveyefficientinferencelarge).
@@
However, LLMs have limitations that stem from their inability to perform deep dynamic analysis. Because they do not execute code (xue2024; yang2024llm), they cannot analyze real-time state changes, such as tracking memory usage during execution or handling complex input/output scenarios dynamically. Their pre-trained nature also makes it challenging to reason about concurrency and parallelism, which involves understanding multi-threaded behavior, race conditions, or potential deadlocks. This type of analysis (sun2024llmruntimeerrorhandler; chen2024reasoningruntimebehaviorprogram; krishna2024codellmdevkitframeworkcontextualizingcode) requires direct code execution and monitoring of parallel processes, which LLMs cannot achieve. Furthermore, LLMs lack the capability to handle real-time data changes or interact with code based on live data, as they primarily operate on static representations and their training knowledge. This limitation prevents them from fully simulating execution environments or adapting to live input changes, highlighting an important gap compared to runtime analysis tools.
Although LLMs are inherently limited in performing extensive dynamic analysis, particularly in complex scenarios involving real-time state changes, concurrency, and live data interactions, our framework demonstrates a path to extend their utility through strategic assistance. By combining static analysis techniques with the logical reasoning (parmar2024; li2024relational) and domain knowledge that LLMs inherently possess, we can bridge this gap effectively.
@@ 2.2. Compiler-Assisted LLM Reasoning: Combining World Knowledge with Static Analysis
When these compiler generated static artifacts are paired with an LLM’s capabilities in reasoning (sun2024llmruntimeerrorhandler; chen2024reasoningruntimebehaviorprogram; krishna2024codellmdevkitframeworkcontextualizingcode) about control flow, algorithmic complexity, and domain-specific behavior, a comprehensive analysis of code behavior emerges. The LLM can use static analysis insights to simulate runtime behaviors, assess interactions, and predict the impact of changes more effectively. For instance, a call graph might reveal complex function dependencies, while the LLM can interpret how these dependencies influence runtime performance or suggest optimizations. However, performing such analysis within the LLM itself can be computationally expensive, shifting focus from estimating hot functions to more generalized codebase analysis. Since it is an interactive tool, LLMs typically budget a certain time interval for getting the response back to the user and within that budget different phases within the LLM must fit. To offload the burden of analysis and to utilize that time for some other inferencing phases, we assist the LLM with logical cues such as the set of all functions used within the benchmark, function call graphs, and some hotness indicators such as whether or not called within loops etc. This approach reduces the model’s search space, enabling more efficient and targeted inference by emphasizing functions that are statistically and structurally significant. This combination of static analysis and probabilistic guidance empowers the LLM to deliver more accurate and efficient code insights, bridging gaps in its inherent capabilities (hadi2024).
@@ 2.3. Alternative Approach: Profile Generalization & Generative Models
In order to utilize this WPP based function profile data for profile generalization and accomplishing profile-driven tasks, we train generative models on WPP based function profile data, which predicts the upcoming tokens (function calls) based on a prefix of tokens, and simultaneously adjusts to other execution scenarios (different inputs). Deep-learning generative models such as Recurrent Neural Networks (RNNs) can estimate the probability of the next token (at timestep $t$) in a sequence, as function of the previous token (at timestep t1t-1) and an internal short-term memory (which propagates information across the sequence). The utility of RNNs stems from the fact that the output token in the previous execution timestep acts as the input token for the current execution timestep, which simplifies the process of querying and inferencing the model. Popular generative models such as GANs (goodfellow2020generative), Diffusion Models (yang2023diffusion) etc, are extremely compute-heavy ($\sim 10^{20}$ FLOPS per epoch), and require extensive resources for training (100s of GPU hours). In contrast, we show that simple RNNs with one hidden layer (vanilla RNNs) are capable of learning the dynamic function call chains (§5.5).
@@ 3.1. Phase 1: Filtering Search Space with Domain Knowledge Deduction
Our initial experiments showed that once the application domain is established, a direct correlation can be established between algorithm-specific cost models, and candidate hotspot functions. For instance, 510.parest_r510.parest\_r from SPEC2k17 (bucek2018spec) implements a finite element solver for biomedical imaging (Listing 1). Based on this domain knowledge, the LLM can immediately deduce that quadrature evaluations, gradient computations, and scatter operations are computationally intensive. This leads to the isolation of s_val, s_grad, and cpy_matrix as potential hotspots, sidestepping the high-level assemble_matrix wrapper. Similarly, 505.mcf_r505.mcf\_r from the same benchmark suite adopts a primal network simplex algorithm to solve the minimum cost flow problem (Listing 2). This deduction enlightened the LLM to attribute runtime to the domain specific constructs such as pricing (cost_compare), sorting (spec_qsort), and implicit-arc maintenance (bea_is_dual_infeasible) functions that dominate execution at scale, instead of the top-level driver primal_net_simplex. We observed a similar pattern in other application domains, such as molecular dynamics (neighbor list construction and electrostatics kernels), fluid mechanics (collide-and-stream sweeps), and image processing (convolution, and resampling), where domain knowledge consistently filtered out high-level drivers in favor of the computational kernels.
@@
Listing 1: Finite Element Solver code snippet from 510.parest_r in SPEC2017 (bucek2018spec). While the wrapper dominates inclusive time, the true hotspots are the inner quadrature and scatter kernels, based on domain knowledge.
@@
Listing 2: Primal Simplex code snippet from solving minimum cost flow in 505.mcf_r in SPEC2017 (bucek2018spec). Despite enclosing the optimization loop, the runtime is dominated by the inner pricing and sorting routines